The level of nitrogen oxides (NOX) in the exhaust of a particular car model varies with mean 0.9 grams per mile and standard deviation 0.2 grams per mile .
(a) What sample size is needed so that the standard deviation of the sampling distribution is 0.01 grams per mile ?

Answer by Theo(13342)

mean = .9

standard deviation = .2

standard error = standard deviation of sampling distribution.

formula for that is s = standard deviation / square root of sample size.

to make s = .01, the formula becomes .01 = .2 / sqrt(n)

n is the sample size.

multiply both sides of this equation by sqrt(n) and divide both sides of this equation by .01 to get:

sqrt(n) = .2 / .01.

solve for sqrt(n) to get sqrt(n) = 20

solve for n to get n = 20^1 = 400.

when n = 400, the standard error will be .2 / sqrt(400) = .2 / 20 = .01.